Wednesday, January 11, 2006

No voltage supplied for Anode Hiptronics power supply

The repaired Lecroy oscilloscope was sent back this morning, we can use it to record the waveforms from now.

I received the new relays this afternoon, the circuit breaker was still kicked off after settling down them. Mike suggested to check the voltage phases. I found that I connected K3-1 and K3-3 incorrectly. The filament voltage became normal(~5.4 V DC, each side 2.7 V) after switching these two connectors. Then I tried to test the whole system, but this time there was no voltage spplied to Hiptronics High Voltage Power Supply. I measured the voltages on the MO5B board, the TB01-01 and TB01-03 both are zero, they are connected to MO5ATB03-01 and MO5ATB02-09 respectively(circuit drawing shown right). So I will check MO5A tomorrow.

No comments: